A power boat has a 4 hour supply of gasoline. How far can this boat travel from the marina if the rate out against the current is 40 km/h and the rate back in with the current is 60 km/h?
Answer by checkley71(8403)

40*X=60*(4-X) 0R 40X=240-60X OR 100X=240 OR X=2.4 HOURS OUT @ 40KM/H AND
4-2.4 OR 1.6 HOURS FOR THE RETURN TRIP. THUS THE TOTAL TRIP IS 2.4 HOPURS @ 40 OR 96 KM AND THE RETURN TRIP IS 60*16 OR 96 KM.